if the area of a triangle is 375 square meters, what is the measure of the height if the base is 25 meters
?
1
Expert's answer
2012-02-16T09:00:06-0500
Area of the triangle S=0.5*h*a where h-height, a -corresponding base
whence height h=2S/a
Substituting given values we obtain
h=2*375/25=30 meters